OpenStudy (anonymous):

A triangle has a perimeter of 161 miles. The length of each o the two smaller sides of the triangle is two-thirds the length of the longest side. Find the length of each side of the triangle.

OpenStudy (anonymous):

OMG I KNOW YOU

OpenStudy (anonymous):

 Solve for X... 2/3x + 2/3x + 3/3x = 7/3x. 7/3x = 161. x = 69. Longest side is 69. Solve for the shortest side. One short side is 2/3 of that. That gives you 46. Your triangle has the longest side of 69 units, and the two shorter sides of 46.

OpenStudy (anonymous):

let three side be x, y, z now smaller sides i.e x, y x=y= 2/3z perimeter of triangle=x+y+z=161 2/3z+2/3z+z=161 7/3z=161 z=69 from z we can calculate x & y i.e 46 three sides are 46, 46,& 69 miles

OpenStudy (anonymous):

I don't understand how you go 7/3?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

u r gr8

OpenStudy (anonymous):

High you are (Branchez Remix) - Whats So Not

OpenStudy (anonymous):

good song